OpenStudy (anonymous):

The equation of a line y=1/4x-2 What is the equation of the line that is parallel to the first line and passes through (4, –2)?

OpenStudy (jdoe0001):

what's the slope of y=1/4x-2 ?

OpenStudy (anonymous):

the slope is 1/4

OpenStudy (da_scienceman):

I will calcilate the gradient or slope of the second line which is parallel to the first as follows: y1.y2=-1, so I feel the seond line has slope -4. The n using (y-y2)=m2(x-x2), I substitue x2=4, y2=-2 to get: y-(-2)=-4(x-4). Simplify to get the required equation of the second line.

OpenStudy (anonymous):

y = –4x + 14 thanks

OpenStudy (anonymous):

No its y=1/4x-3, the slope would be -4 if it was a perpendicular line

OpenStudy (anonymous):

oh i suck so bad at this thank you omfg

OpenStudy (anonymous):

Just remember that whenever its parallel, it has the same slope, so you just need to plug in your x and y values to figure out the y-intercept

OpenStudy (da_scienceman):

ok right the slopes are same, so it will be (y-(-2))=1/4(x-4)...sorry. I thought it was perpendicular! Thx for pointing that out!

OpenStudy (da_scienceman):

so it will be 4y=x-12! sorry